Show whether the following equation are exact and hence solve the equation.


a) 2x(ye^x² -1)dx + e^x²dy= 0


b) ( 6x⁵y³ + 4x³y⁵) dx + (2x⁶y² - 5x⁴y⁴) dy =0

1
Expert's answer
2022-05-03T06:44:23-0400

"a) 2x(ye^{x^2}-1)dx+e^{x^2}dy =0"

"M(x,y)=2x(ye^{x^2}-1)" and "N(x,y)=e^{x^2}"

"M_y=2xe^{x^2}=N_x"

The given differential equation is exact.

"\int M(x,y)dx=ye^{x^2}-x^2+g(y)"

"\int N(x,y)dy =ye^{x^2}+h(x)"

Answer: "f(x,y)=ye^{x^2}-x^2+C."


"b) (6x^5y^3+4x^3y^5)dx+(2x^6y^2-5x^4y^4)dy=0"

"M(x,y)=6x^5y^3+4x^3y^5" and "N(x,y)=2x^6y^2-5x^4y^4"

"M_y=18x^5y^2+20x^3y^4" and "N_x(x,y)=12x^5y^2-20x^3y^4"

"M_y\neq N_x"

The given differential equation is not exact.